# Distinguishing Neighborhood Representations Through Reverse Process of GNNs for Heterophilic Graphs

### v1

@@ Abstract
Graph Neural Network (GNN) resembles the diffusion process, leading to the over-smoothing of learned representations when stacking many layers. Hence, the reverse process of message passing can sharpen the node representations by inverting the forward message propagation. The sharpened representations can help us to better distinguish neighboring nodes with different labels, such as in heterophilic graphs. In this work, we apply the design principle of the reverse process to the three variants of the GNNs. Through the experiments on heterophilic graph data, where adjacent nodes need to have different representations for successful classification, we show that the reverse process significantly improves the prediction performance in many cases. Additional analysis reveals that the reverse mechanism can mitigate the over-smoothing over hundreds of layers.
@@ 1 Introduction
Graph neural networks (GNN) have emerged as an important tool for learning relational data. Earlier attempts aim to learn the node representations from graphs based on a message-passing mechanism. The message-passing neural network framework shows partial success with the homophilic graphs, where the nodes with the same labels are likely to be connected. When the heterophilic graphs, where node labels significantly differ from those of their neighbors, are considered, the models based on the homophilic assumption often perform worse than the naive neural network architectures without considering the relationship between nodes (Zhu et al., 2020).
Two classes of GNNs have been proposed to overcome the limitation on heterophilic graphs. The first class aims to aggregate the information selectively from the neighbors (Chien et al., 2020; Li et al., 2022). For example, aggregating relevant information from multi-hop neighbors or relevant features from neighbors has been proposed in this direction (Zhu et al., 2020; Liang et al., 2023). On the other hand, the other class of methods highlights the importance of initial node features before aggregation. Adding skip-connection to the initial node features or isolating the original node features for prediction has been proposed in this direction (Zhu et al., 2020; Song et al., 2023).
@@
Figure 1: Visualized node representations over the forward and reverse processes in Minesweeper. Top: our approach with the forward and reverse processes. Bottom: a classical GCN with a forward process only. The original node features are smoothed over the forward process, whereas the features are sharpened over the reverse process. Visualization details are provided in Section 4.1.3.
Aggregation is a fundamental mechanism of all GNN models, and heterophilic GNNs are no exception. With a very clever technique, one can mitigate the propagation of irrelevant information for the first few layers. However, GRAND (Chamberlain et al., 2021) shows that GNN can be seen as a discretization of a heat diffusion equation. The diffusion perspective implies that the node representations will eventually reach the stationary state, with which the nodes cannot be distinguished. Therefore, if we cannot isolate all irrelevant nodes in the aggregation process, which is typically impossible, the GNN will eventually aggregate information from irrelevant nodes as the model propagates over and over.
@@
In this work, we claim not to forcefully correct the diffusive nature of the GNNs. Instead, we propose to use the reverse process of the aggregation. The aggregation process is known to smooth the node representations; hence, its reverse process can sharpen the node representations and make the neighborhood representations more distinguishable. To illustrate our intuition, we showcase our experimental results on the Minesweeper dataset, a well-known heterophilic dataset, in Figure 1. In Minesweeper, a board is a grid-structured graph where each node is initialized with the number of mines in the adjacent nodes $\mathbf{X}^{(0)}$, and the goal is to classify the location of mine $\mathbf{Y}$ correctly. The bottom row visualizes the learned node representations with the GCN (Kipf & Welling, 2017), and the top row visualizes the representations with our approach. With the forward-only method, such as GCN, the learned representations cannot be sharpened enough to distinguish the target nodes from the others, whereas when the reverse process is applied to the initial features, we can obtain a sharpened representation from the backward process.
@@
The experimental results on heterophilic datasets show that the reverse process improves the prediction performance compared with the forward-only models. Our approach produces sharper representations and successfully stacks deep layers to capture long-range dependencies, which are crucial for performance on heterophilic datasets. The experiments on homophilic datasets confirm that the reverse process does not harm the prediction performance when the aggregation mechanism is sufficient. Our investigation reveals that with the reverse process, one can stack hundreds, even a thousand layers, without suffering over-smoothing. Additional experiments with the low-rate label datasets confirm that the reverse process design is effective for tasks where the long-range interaction between nodes is inevitable.
@@ 2 Related Work
Most studies on heterophilic data focus on identifying nodes with similar characteristics even among non-adjacent ones for aggregation. GPR-GNN (Chien et al., 2020) utilizes a trainable generalized PageRank for feature aggregation, learning important neighborhood ranges from the data and emphasizing information within those ranges for aggregation. CPGNN (Zhu et al., 2021) introduces a learnable compatibility matrix to capture the information of non-adjacent homophilic nodes. FAGCN (Bo et al., 2021) introduces a GNN framework with a self-gating mechanism to adaptively use low-frequency and high-frequency signals. FSGNN (Maurya et al., 2022) proposes soft feature selection, wherein it adaptively selects node features with different properties. GloGNN (Li et al., 2022) employs a coefficient matrix that represents node-to-node relationships for aggregation, allowing the aggregation of information from all nodes. GBK-GNN (Du et al., 2022) proposes a bi-kernel graph neural network that separately handles homophilic and heterophilic nodes. It uses a selection gate to predict whether a node is homophilic or heterophilic and obtains features using the corresponding kernel based on the prediction. LRGNN (Liang et al., 2023) uses a low-rank approximation to compute a label relationship matrix, employing it for signed message passing.
However, aggregation still causes global node representations to become similar, known as over-smoothing, leading to performance degradation on heterophilic datasets. To overcome this issue, H2GCN (Zhu et al., 2020) and OGNN (Song et al., 2023) proposes to preserve non-aggregated representation separately. H2GCN learns node representation by separating ego-embedding and neighbor-embedding and employing intermediate representations. OGNN proposes ordering message passing to prevent the mixing of messages from different hops.
@@ 3.1 Framework
In contrast to a typical GNN layer, we propose a GNN layer $g$ that performs the opposite role by reversing diffusion and re-concentrating diffused information. Our main idea is to design an inverse function of a message-passing GNN layer $f$, and $g=f^{-1}$ to perform the reversion. Due to the diffusive nature of the GNN layer $f$, its inverse form $g$ is expected to have two properties: 1) $g$ cancels the smoothing effect, producing sharper representations and thus leading to 2) stacking multiple layers of $g$ that do not suffer from the over-smoothing issue.
@@ 3.2 Reverse Diffusion Based on GRAND
Within the framework of diffusion, the time parameter serves as a continuous layer, similar to the concept used in NeuralODEs (Chen et al., 2018). Using Equation 1, GRAND produce node representations at time $T_{D}>0$ by:
@@
where numerical techniques like the Euler method are used to solve integration. Since Equation 1 models the property of heat reaching equilibrium over time, the node representations obtained through Equation 2 become diffused as time progresses. Conversely, tracing back in time allows us to observe the concentrated form of heat before diffusion. Utilizing Equation 1, past node representations at time $T_{R}<0$ can be calculated as follows:
@@
Equation 3 reverses the diffusion process, enabling us to obtain sharpened representations.
@@ 3.3 Reverse Process Based on GNN with residual connections
We have explored the design of reverse process in widely used two message-passing GNN structures: graph convolutional network (GCN) (Kipf & Welling, 2017) and graph attention network (GAT) (Veličković et al., 2018). Both GCN and GAT model a node representation through 1) an aggregation step, where neighborhood representations are combined, and 2) an update step, where the aggregated representations are merged into the target node representation. Let $\hat{\mathbf{A}}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ encode neighborhood structure in a graph, and $\mathbf{W}\in\mathbb{R}^{d\times d}$ be a matrix of learnable parameters. With the application of skip-connections (He et al., 2016), the GCN and GAT layers can be formalized as
$f(\mathbf{X}^{(\ell)})=\mathbf{X}^{(\ell+1)}$ $=\mathbf{X}^{(\ell)}+\sigma(\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W})$
$=\mathbf{X}^{(\ell)}+h(\mathbf{X}^{(\ell)}),\;$
where $\sigma(\cdot)$ represents a non-linear activation function. $\hat{\mathbf{A}}$ is the renormalized adjacency matrix in GCN and a learnable attention matrix in GAT.
According to Behrmann et al. (2019), the inverse of the GNN layer $f^{-1}$ exists if the $\text{Lip}(h)<1$, where $\text{Lip}(h)$ is Lipschitz constant of $h$. In this case, $f^{-1}(\mathbf{X}^{(\ell)})$ can be computed via fixed point iteration as described in Algorithm 1.
@@
  Input: output of residual layer $\mathbf{X}^{(\ell)}$, residual block $h$, number of fixed-point iterations $M$
@@
If $h$ is a simple feed-forward network, we can constrain its Lipschitz constant by normalizing the weight matrix through the spectral norm. However, computing the spectral norm is time-consuming. Instead, we use an upper bound of $\text{Lip}(h)$ in our case. An upper bound of $\text{Lip}(h)$ is calculated as follows since $\lVert{\mathbf{XY}}\rVert_{F}\leq\lVert{\mathbf{X}}\rVert_{F}\lVert{\mathbf{Y}}\rVert_{F}$:
$\text{Lip}(h)=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\leq\lVert{\hat{\mathbf{A}}}\rVert_{2}\lVert{\mathbf{W}}\rVert_{F},$ (5)
where $\lVert{\cdot}\rVert_{F}$ denotes Frobenius norm, and $\lVert{\cdot}\rVert_{2}$ spectral norm. In GCN, $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix $\mathbf{A}$ with added self-loops and $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$. A spectral norm of normalized adjacency matrix $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ in GCN structure. Also, $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ when $\hat{\mathbf{A}}$ is right-stochastic, which is the case of GAT. Unlike Behrmann et al. (2019), where power iteration is performed to compute the Lipschitz constant, in GCN and GAT, it is sufficient to renormalize the weight matrix through the Frobenius norm of it. The upper bound normalization reduces the time complexity at the expense of the exact Lipschitz constant calculation. In experiments, we find that the Frobenius upper bound can still result in a competitive performance.
When the scaling coefficient $c<1$ is given, $\mathbf{W}$ is normalized to $\frac{c\mathbf{W}}{\lVert{\mathbf{W}}\rVert_{F}}$ if $c<\lVert{\mathbf{W}}\rVert_{F}$, in order to satisfy $\text{Lip}(h)<c$. When multi-head attention with $K$ heads is employed for GAT, we use the average of representations obtained from multiple heads as a node representation. Specifically, parameters of $k$-th head $\mathbf{W}^{(k)}$ is normalized to
$\mathbf{W}^{(k)}\leftarrow\frac{c\mathbf{W}{(k)}}{\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}},\;\forall k,\;$ (6)
since the Liptschitz upper bound result in $\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$. The derivation of Liptschitz upper bound for multi-head attention is provided in Appendix C.
While the time complexity of a GCN mainly depends on the number of the forward propagation $L_{D}$, the complexity of the reverse process depends on the number of fixed point iterations $M$ and of the number of reverse layers $L_{R}$. In our implementation, we run the fixed point iteration until convergence and backpropagate over the iterations. We provide the time and memory complexity analysis in Table 1, and the proof is provided in Appendix A. An analysis on $M$ and the run-time with varying $L_{R}$ in real experiments are provided in Section 4.3.
@@
Forward Time $O(L_{D}|\mathcal{E}|d+L_{D}|\mathcal{V}|d^{2})$ $O((L_{D}+M^{2}L_{R})|\mathcal{E}|d+(L_{D}+M^{2}L_{R})|\mathcal{V}|d^{2})$
Forward Memory $O(|\mathcal{E}|+L_{D}|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+(L_{D}+L_{R})|\mathcal{V}|d+d^{2})$
Backward Time $O(L_{D}^{2}|\mathcal{E}|d+L_{D}^{2}|\mathcal{V}|d^{2})$ $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$
Backward Memory $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$ $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$
Table 1: Space and time complexity of GCN for the forward and backward processes. $L_{D}$ and $L_{R}$ represent the number of forward and reverse steps, respectively, $M$ is the number of fixed point iterations, and $d$ is the dimensionality of the node representation.

### v2

# Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs

@@ Abstract
Graph Neural Network (GNN) resembles the diffusion process, leading to the over-smoothing of learned representations when stacking many layers. Hence, the reverse process of message passing can produce the distinguishable node representations by inverting the forward message propagation. The distinguishable representations can help us to better classify neighboring nodes with different labels, such as in heterophilic graphs. In this work, we apply the design principle of the reverse process to the three variants of the GNNs. Through the experiments on heterophilic graph data, where adjacent nodes need to have different representations for successful classification, we show that the reverse process significantly improves the prediction performance in many cases. Additional analysis reveals that the reverse mechanism can mitigate the over-smoothing over hundreds of layers. Our code is available at https://github.com/ml-postech/reverse-gnn.
@@ 1 Introduction
Graph neural networks (GNNs) have emerged as an important tool for learning relational data. Earlier attempts aim to learn the node representations from graphs based on a message-passing mechanism. The message-passing neural network framework shows partial success with the homophilic graphs, where the nodes with the same labels are likely to be connected. When the heterophilic graphs, where node labels significantly differ from those of their neighbors, are considered, the models based on the homophilic assumption (McPherson et al., 2001) often perform worse than the naive neural network architectures without considering the relationship between nodes (Zhu et al., 2020).
@@
Figure 1: Visualized node representations over the forward and reverse processes in Minesweeper. Top: our approach with the forward and reverse processes. Bottom: a classical GCN with a forward process only. The original node features are smoothed over the forward process, whereas the features are more distinguishable over the reverse process. Visualization details are provided in Section 4.1.3.
To learn the node representations of heterophilic graphs, a GNN needs to capture long-range interactions between nodes, leading to the stacking of multiple message-passing layers (Li et al., 2022; Rusch et al., 2023). However, many studies empirically and theoretically identify that GNN tends to smooth the node representations over the layers, and eventually, the learned representations are likely to be similar, known as the over-smoothing issue (Chen et al., 2020; Rusch et al., 2023). Furthermore, GRAND (Chamberlain et al., 2021) shows that GNN can be seen as a discretization of a heat diffusion equation. The diffusion perspective implies that learning distinguishable node representation with deep GNN is challenging since heat only diffuses to reach equilibrium, where node representation becomes indistinguishable.
In this work, we claim not to forcefully correct the diffusive nature of the GNNs. Instead, we propose to use the reverse process of the aggregation. The aggregation process is known to make the node representations similar; hence, its reverse process can make the neighborhood representations more distinguishable. Revisiting the diffusion perspective, applying the reverse process means learning the states in the past, which are further away from equilibrium and more distinguishable.
@@
To illustrate our intuition, we showcase our experimental results on the Minesweeper dataset, a well-known heterophilic dataset (Platonov et al., 2023b), in Figure 1. In Minesweeper, a board is a grid-structured graph where each node is initialized with the number of mines in the adjacent nodes $\mathbf{X}^{(0)}$, and the goal is to classify the location of mine $\mathbf{Y}$ correctly. The top row visualizes the learned node representations with our approach, and the bottom row visualizes the representations with the GCN (Kipf & Welling, 2017). With the forward-only method, such as GCN, the learned representations often fail to obtain distinguishable representations for classification. However, when the reverse process is applied to the initial features, we can obtain a distinguishable representation from the backward process.
@@
The experimental results on heterophilic datasets show that the reverse process improves the prediction performance compared with the forward-only models. Our investigation reveals that the reverse process produces distinguishable representations and enables the stacking of hundreds, even a thousand layers, mitigating over-smoothing. Successfully stacking deep layers allows for the capture of long-range dependencies, which are crucial for performance on heterophilic datasets. The experiments on homophilic datasets confirm that the reverse process does not harm the prediction performance when the aggregation mechanism is sufficient.
@@ 2 Related Work
Most studies on heterophilic data focus on identifying nodes with similar characteristics even among non-adjacent ones for aggregation. GPR-GNN (Chien et al., 2020) utilizes a trainable generalized PageRank for feature aggregation, learning important neighborhood ranges from the data and emphasizing information within those ranges for aggregation. CPGNN (Zhu et al., 2021) introduces a learnable compatibility matrix to capture the information of non-adjacent homophilic nodes. FSGNN (Maurya et al., 2022) proposes soft feature selection, wherein it adaptively selects neighbors to aggregate with different hop distances. GloGNN (Li et al., 2022) employs a coefficient matrix that represents node-to-node relationships for aggregation, allowing the aggregation of information from all nodes. GBK-GNN (Du et al., 2022) proposes a bi-kernel graph neural network that separately handles homophilic and heterophilic nodes. It uses a selection gate to predict whether a node is homophilic or heterophilic and obtains features using the corresponding kernel based on the prediction. LRGNN (Liang et al., 2023) uses a low-rank approximation to compute a label relationship matrix, employing it for signed message passing.
However, aggregation still causes global node representations to become similar, known as over-smoothing, leading to performance degradation on heterophilic datasets. To overcome this issue, H2GCN (Zhu et al., 2020) and Ordered GNN (Song et al., 2023) proposes to preserve non-aggregated representation separately. H2GCN learns node representation by separating ego-embedding and neighbor-embedding and employing intermediate representations. Ordered GNN proposes ordering message passing to prevent the mixing of messages from different hops.
Several studies propose methods to adaptively learn appropriate filters that can handle various graph structures. FAGCN (Bo et al., 2021) introduces a GNN framework with a self-gating mechanism to adaptively use low-frequency and high-frequency signals. JacobiConv (Wang & Zhang, 2022) uses Jacobi bases for spectral filter, whose orthogonality and flexibility enable adaptation to a wide range of graph signal densities. ACM-GCN (Luan et al., 2022) utilizes a filterbank which combines low-pass and high-pass filters, and adaptively considers node-wise local information.
@@ 3.1 Framework
In contrast to a typical GNN layer, we propose a GNN layer $g$ that performs the opposite role by reversing diffusion and re-concentrating diffused information. Our main idea is to design an inverse function of a message-passing GNN layer $f$, with $g=f^{-1}$ to perform the reversion. Due to the diffusive nature of the GNN layer $f$, its inverse form $g$ is expected to have two properties: 1) $g$ cancels the smoothing effect, producing distinguishable representations that mitigate the over-smoothing issue and thus 2) leading to stacking multiple layers of $g$.
@@ 3.2 Reverse Diffusion Based on GRAND
Within the framework of diffusion, the time parameter serves as a continuous layer, similar to the concept used in NeuralODEs (Chen et al., 2018). Using Equation 1, GRAND produce node representations at time $T_{F}>0$ by:
@@
where numerical techniques like the Euler method are used to solve integration. Since Equation 1 models the property of heat reaching equilibrium over time, the node representations obtained through Equation 2 become diffused as time progresses. Conversely, tracing back in time allows us to observe the concentrated form of heat before diffusion. Utilizing Equation 1, node representations at a past time $T_{R}<0$ can be calculated as follows:
@@
Equation 3 reverses the diffusion process, enabling us to obtain distinguishable representations.
@@ 3.3 Reverse Process Based on GNN with Residual Connections
We have explored the design of reverse process in widely used two message-passing GNN structures: graph convolutional network (GCN) (Kipf & Welling, 2017) and graph attention network (GAT) (Veličković et al., 2018). Both GCN and GAT model a node representation through an aggregation step, where neighborhood representations are combined, and an update step, where the aggregated representations are merged into the target node representation. Let $\hat{\mathbf{A}}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ encode neighborhood structure in a graph, and $\mathbf{W}\in\mathbb{R}^{d\times d}$ be a matrix of learnable parameters. With the application of skip-connections (He et al., 2016), the GCN and GAT layers can be formalized as
$f(\mathbf{X}^{(\ell)})=\mathbf{X}^{(\ell+1)}$ $=\mathbf{X}^{(\ell)}+\sigma(\hat{\mathbf{A}}\mathbf{X}^{(\ell)}\mathbf{W})$ (5)
$=\mathbf{X}^{(\ell)}+h(\mathbf{X}^{(\ell)}),\;$ (6)
where $\sigma(\cdot)$ represents a non-linear activation function. $\hat{\mathbf{A}}$ is the renormalized adjacency matrix in GCN and a learnable attention matrix in GAT. We note that when using multi-head attention, we take averaging approach as in GRAND to keep invertibility.
According to Behrmann et al. (2019), the inverse of the GNN layer $g=f^{-1}$ exists if the $\text{Lip}(h)<1$, where $\text{Lip}(h)$ is Lipschitz constant of $h$. With the contractive nonlinear activations like ReLU, ELU, and tanh, $\text{Lip}(h)<1$ is satisfied if
$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}<1\;,$ (7)
where $\lVert{\cdot}\rVert_{F}$ denotes Frobenius norm. When the condition is guaranteed, $g(\mathbf{X}^{(\ell)})$ can be computed via fixed point iteration as described in Algorithm 1, resulting in $\mathbf{X}^{(\ell-1)}$.
To ensure the invertibility of $f$ throughout the entire training procedure, we enforce the weight matrix $\mathbf{W}$ to satisfy the condition. Since it is difficult to optimize the weight matrix while satisfying the condition, we normalize the weight matrix after each gradient descent step. Specifically, given that the left side of Equation 7 is upper bounded by
$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}\mathbf{XW}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\leq\lVert{\hat{\mathbf{A}}}\rVert_{2}\lVert{\mathbf{W}}\rVert_{F}\;,$ (8)
where $\lVert{\cdot}\rVert_{2}$ denotes spectral norm, we use the upper bound to normalize the weight matrix.
Note that the spectral norm of $\hat{\mathbf{A}}$ is straightforward for the two models that we considered as baselines: GCN and GAT. In GCN, $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix $\mathbf{A}$ with added self-loops and $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$. A spectral norm of renormalized adjacency matrix $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$. In GAT, $\lVert{\hat{\mathbf{A}}}\rVert_{2}=1$ since $\hat{\mathbf{A}}$ is right-stochastic. Therefore, normalizing the weight matrix through its Frobenius norm is sufficient to guarantee the condition in Equation 7. The upper bound normalization reduces the time complexity at the expense of the exact supremum calculation. In experiments, we find that the Frobenius upper bound can still result in a competitive performance.
@@
  Input: output of residual layer $\mathbf{X}^{(\ell)}$, residual block $h$, the number of fixed-point iterations $M$
@@
When the scaling coefficient $c<1$ is given, $\mathbf{W}$ is normalized to $\frac{c\mathbf{W}}{\lVert{\mathbf{W}}\rVert_{F}}$ if $c<\lVert{\mathbf{W}}\rVert_{F}$, in order to satisfy $\text{Lip}(h)<c$. When multi-head attention with $K$ heads is employed for GAT, parameters of $k$-th head $\mathbf{W}^{(k)}$ is normalized to $\frac{c\mathbf{W}^{(k)}}{\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}}$ for all $k$, since the upper bound result in $\frac{1}{K}\sum_{k=1}^{K}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$. The derivation of the upper bound for multi-head attention is provided in Appendix C. Since residual block $h$ is an operator on a Banach space, and we constraint the $\text{Lip}(h)<1$, the convergence of Algorithm 1 is guaranteed by the Banach fixed point theorem (Behrmann et al., 2019). Inversion error in practice are reported in Section 4.2.
While the time complexity of a GCN mainly depends on the number of the forward layers $L_{F}$, the complexity of the reverse process depends on the number of fixed point iterations $M$ and of the number of reverse layers $L_{R}$. In our implementation, we run the fixed point iteration until convergence and backpropagate over the iterations. We provide the time and memory complexity analysis in Table 1, and the proof is provided in Appendix A. An analysis on $M$ and the run-time with varying $L_{R}$ in real experiments are provided in Section 4.3.
@@
Forward Time $O(L_{F}|\mathcal{E}|d+L_{F}|\mathcal{V}|d^{2})$ $+\;O(M^{2}L_{R}|\mathcal{E}|d+ML_{R}|\mathcal{V}|d^{2}+M^{2}L_{R}d^{3})$
Forward Memory $O(|\mathcal{E}|+L_{F}|\mathcal{V}|d+d^{2})$ $+\;O(L_{R}|\mathcal{V}|d)$
Backward Time $O(L_{F}^{2}|\mathcal{E}|d+L_{F}|\mathcal{V}|d^{2}+L_{F}^{2}d^{3})$ $+\;O(M^{2}L_{R}^{3}|\mathcal{E}|d+ML_{R}^{2}|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$
Backward Memory $O(|\mathcal{E}|+|\mathcal{V}|d+d^{2})$
Table 1: Space and time complexity of GCN for the forward and reverse processes. We show an additional complexity when using reverse process, for simplicity. ✗ denotes there are no additional complexity. $L_{F}$ and $L_{R}$ represent the number of forward and reverse layers, respectively, $M$ is the number of fixed point iterations, and $d$ is the dimensionality of the node representation.